OpenStudy (anonymous):

Can someone show me the steps for: (sin30 + tan45 - csc60) / (sec30 + cos60 + cot45)

OpenStudy (anonymous):

Do you know any of the standard trigonometric ratios?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So then, which part is causing the trouble?

OpenStudy (anonymous):

It usually wouldn't, lol, but I'm having a brain block. Could you please show me the steps to this? I know the answer.

OpenStudy (anonymous):

Well, sin 30 is a half...

OpenStudy (anonymous):

I suppose you could try and simplify it first before actually calculating anything...

OpenStudy (anonymous):

Which way would you like me to do it?

OpenStudy (anonymous):

Whichever way you want to do it. I just want to see how we do this.

OpenStudy (anonymous):

(1/2 + 1 -(2/ sqrt 3)) / ((2/ sqrt 3) + 1/2 +1) = -1?

OpenStudy (anonymous):

See, that's what I thought too, but the answer in the back of my book is different. The answer there is: \[(43-24\sqrt{3}) / 11\]

OpenStudy (anonymous):

sin 30 is 0.5 tan 45 is 1

OpenStudy (anonymous):

\[\frac{\frac{1}{2} + 1 -\frac{2}{\sqrt 3}}{\frac{2}{\sqrt 3} + \frac{1}{2} +1}=\frac{3\sqrt{3}-4}{3\sqrt{3}+4}\cdot \frac{3\sqrt{3}-4}{3\sqrt{3}-4} = \frac{(3\sqrt{3}-4)^2}{11}=\ldots\]

OpenStudy (anonymous):

Of course, I just reaslized that I can't cancel......duh! (hit self over head).